Which of the following equations can be formed starting with x = 0? (a) 2x + 1 = – 1 (b) x/ 2 + 5 = 7 (c) 3x – 1 = – 1 (d) 3x – 1 = 1
 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
Consider the equation 3x–1 =–1.
Transpose -1 to right hand side then it becomes 1

3x = -1 + 1
x = 0/3
x =0

Thanks